the cost of fuel for running a bus is proportional to the square of the speed generated in km/hr. it costs Rs.48 per hour when the bus is mooving at the speed of 20km/hr .what is the most economical speed if the fixed charges are Rs.108 for an hour,over and above the running charges.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Let cost of fuel be ‘P’ Rs/hr & spped of vehicle be ‘v’ in km/hr.
P \alpha v^{2}
P = av^{2}, a = constant
48 = a(20)^{2}
a = \frac{48}{20^{2}} = \frac{3}{25}
P = \frac{3}{25}v^{2}
There is something missing in the question. How much distance bus need to cover should mention. Otherwise, at
P = 108 (\frac{Rs.}{hr})
108 = \frac{3}{25}v^{2}
v = 30 (\frac{km}{hr})
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty